# Prior-based Noisy Text Data Filtering: Fast and Strong Alternative For Perplexity

### v2

@@ B.1 Preserving minority data
As data filtering removes noisy data, which is often found in outliers, it inherently risks discarding valuable minority data (e.g., a highly technical paper with rare terminology, a corpus including minority languages) as a trade-off. Minimizing this trade-off is the general goal of research on filtering.
@@
Table 2 shows that the no-filter setting achieves the highest symbolic understanding, while the PPL-based filter shows very low symbolic understanding, even with the improved overall score. This clearly illustrates the trade-off: as filtering is applied, minority languages are pruned away. In contrast, the prior-based filter shows much higher symbolic performance compared to the PPL-based filter, while also maintaining a higher overall score. This indicates a significantly better trade-off of our method. We provide additional evidence in §B.1.2.
@@
B.1.1 Additional methods to address trade-off
@@
B.1.2 Assessing data loss of prior-based filter in controlled setting
@@
B.2 Sensitivity
Table 6: Overlaps of outliers between the tokenizers of each model and GPT-2.
$e$ LLaMA-3-8B T5-small
0.01 0.7734 0.8978
0.1 0.7558 0.8709
0.2 0.7332 0.8607
0.5 0.7349 0.8588
Sensitivity to tokenizer We additionally compare two tokenizers: LLaMA-3-8B (UTF-8-based, vocab size 128K), and T5-small (SentencePiece-based, vocab size 32K). GPT-2 tokenizer (mainly used in our paper) has a vocab size of 50K. We examine the overlap of outliers (top and bottom $e/2\%$). The outliers appear to be largely consistent across tokenizers, with the T5 exhibiting slightly better alignment than LLaMA-3 (Table 6).
We further conducted an experiment to determine which tokenizer yields better performance, and the result is reported in Table 4. Filtering based on GPT-2 and LLaMA-3 tokenizers performs almost identically, while T5 shows performance degradation. This can be interpreted as follows: We propose three possible interpretations: (1) There may exist a threshold of optimal granularity, which lies between 32K and 50K. (2) Using the tokenizer paired with the model may offer stability. Nevertheless, as demonstrated, priors can be computed even with tokenizers mismatched to the model. If an optimal tokenizer exists, it can be freely adopted at any time.
@@ B.2 Sensitivity
Regarding the threshold ratio of outliers, we follow the optimal threshold of the PPL-based filter. [4] extensively evaluated PPL-based filtering across different selection rates (25%, 50%, 75%) and concluded that 50% was the most effective. Nevertheless, we explore an additional approach for verification (“elbow” in Table 4). As illustrated in 2(b), we normalize $\mu$, $\sigma$, rank to a 010-1 scale, and define boundaries where the gradient crosses 1-1 (on both sides), resulting in an 82% threshold of the data. However, the overall performance was lower than the 50% threshold, which is consistent with the observation of [4].
B.3 Correlation between prior and PPL
@@
Appendix C Details on experiments
C.1 Scores for each benchmark
@@
ARC easy
BIG-bench
wikidata
TriviaQA COPA OBQA PIQA HellaSwag LAMBADA
Winograd
Winogrande
@@
BIG-bench
algorithms
BIG-bench
dyck lan-
guages
BIG-bench
elementary
math QA
BIG-bench
operators
GSM8K SVAMP LSAT-LR LSAT-RC SAT-English CoQA
@@
C.2 Explanation for experimental settings
C.2.1 Training token duration
We pretrain our baselines for 6B token duration, repeating 2 times on 3B corpus. This is to inherit the plausibility of the setting in [4]. As shown in Figure 1 of [4], the performance ranking of the baselines observed at 5B tokens remains consistent and stable across later stages of training. This shows that performance gaps between baselines are largely established around 5B token duration. Our work leverages this observation.
@@ C.2.1 Training token duration
To elaborate, Section 3.5 of [4] explains that all models are trained with 2× data repetition. Accordingly, the 5B token duration in Figure 1 corresponds to 2.5B tokens of data repeated twice. Our experimental setup uses 3B tokens repeated twice, which makes it directly comparable to the 5B token point in Figure 1 of [4]. Therefore, our experiment reproduces the training regime up to the 5B token mark in [4].
Additionally, [22] also demonstrates similar patterns, where the performance ranking of the baselines remains consistent from the early stage.
Appendix D Details on Benchmarks
@@
Jha et al. [15] also use the MosaicML evaluation gauntlet to perform evaluations in their work. As such, with explicit permission from the authors, we reproduce their text describing the tasks and task categories in the evaluation gauntlet. The following is from Section D of their paper:
@@
  • ARC easy: 2,376 easy four-choice multiple choice science questions drawn from grade 3-9 science exams. [11]
  • BIG-bench wikidata: 20,321 questions regarding factual information pulled from Wikipedia. [40]
  • TriviaQA: 3,000 question-answering dataset; clipped all answers to be at most 10 tokens long to improve speed. [17]
@@
  • COPA: 100 cause/effect multiple choice questions. [35]
  • OBQA (OpenBook QA): 500 four-choice multiple choice questions that rely on basic physical and scientific intuition about common objects and entities. [24]
  • PIQA: 1,838 commonsense physical intuition 2-choice multiple choice questions. [9]
@@
  • HellaSwag: 10,042 multiple choice scenarios in which the model is prompted with a scenario and choose the most likely conclusion to the scenario from four possible options. [45]
  • LAMBADA: 6,153 passages take from books - we use the formatting adopted by OpenAI’s version. [27]
  • Winograd Schema Challenge: 273 scenarios in which the model must use semantics to correctly resolve the anaphora in a sentence. [21]
  • Winogrande: 1,267 scenarios in which two possible beginnings of a sentence are presented along with a single ending. [37]
@@
  • BIG-bench algorithms: 1,320 multiple choice questions. [40]
  • BIG-bench dyck languages: 1000 complete-the-sequence questions. [40]
  • BIG-bench elementary math QA: 38,160 four-choice multiple choice arithmetic word problems. [40]
  • BIG-bench operators: 210 questions involving mathematical operators. [40]
  • GSM8K: 1,319 short, free-response grade school-level arithmetic word problems with simple numerical solutions. [12]
  • SVAMP: 300 short, free-response grade school-level arithmetic word problems with simple numerical solutions. [28]
@@
  • LSAT-LR: 510 passage-based four choice multiple choice questions. [46]
  • LSAT-RC: 268 passage-based four choice multiple choice questions. [46]
  • SAT-English: 206 passage-based four choice multiple choice questions. [46]
  • CoQA: 7,983 passage-based short free response questions. [34]
Appendix E Usage of AI assistants
@@
Appendix F Filtered cases
@@
F.1 Prior stds($\sigma_{\mathtt{d}}\$) based
@@
d
F.2 Prior mean($\mu_{\mathtt{d}}\$)-based

### v4

@@ Appendix D Preserving minority data
As data filtering removes noisy data, which is often found in outliers, it inherently risks discarding valuable minority data (e.g., a highly technical paper with rare terminology, a corpus including minority languages) as a trade-off. Minimizing this trade-off is the general goal of research on filtering. In this section, we discuss our prior-based filtering method from the view of this trade-off.
@@
Table 2 shows that the no-filter setting achieves the highest symbolic understanding, while the PPL-based filter shows very low symbolic understanding, even with the improved overall score. This clearly illustrates the trade-off: as filtering is applied, minority languages are pruned away. In contrast, the prior-based filter shows much higher symbolic performance compared to the PPL-based filter, while also maintaining a higher overall score. This indicates a significantly better trade-off of our method. We provide additional evidence in §D.2.
@@
D.1 Additional methods to address trade-off
@@
Table 6: Performance among baselines pre-trained on Dolma.
Criteria Source Average normalized accuracy World knowledge Commonsense reasoning Language understanding Symbolic problem solving Reading compre- hension
Large (1.5B) model
prior-based mean Dolma 8.50 9.12 10.25 7.45 11.38 4.28
" stds " 8.70 7.28 10.57 9.34 12.40 3.89
" mean + stds " 9.20 9.53 11.27 10.31 11.13 3.79
" " Dolma + Github 9.48 11.47 10.83 8.97 12.22 3.78
D.2 Assessing data loss of prior-based filter in controlled setting
@@
Appendix E Additional analysis
E.1 Sensitivity
@@
Regarding the threshold ratio of outliers, we follow the optimal threshold of the PPL-based filter. Ankner et al. (2024) extensively evaluated PPL-based filtering across different selection rates (25%, 50%, 75%) and concluded that 50% was the most effective. Nevertheless, we explore an additional approach for verification (“elbow” in Table 9). As illustrated in 2(b), we normalize $\mu$, $\sigma$, rank to a 010-1 scale, and define boundaries where the gradient crosses 1-1 (on both sides), resulting in an 82% threshold of the data. However, the overall performance was lower than the 50% threshold. We also present an experiment with a threshold of 25%, which also showed clearly lower performance. These observations were consistent with the observation of Ankner et al. (2024).
Table 9: Performance of large (1.5B) models on Dolma, under different threshold settings.
Threshold Average normalized accuracy World knowledge Commonsense reasoning Language understanding Symbolic problem solving Reading compre- hension
25% 6.21 7.03 7.06 5.99 7.36 3.57
50% 9.20 9.53 11.27 10.31 11.13 3.79
81% (elbow) 8.79 10.04 7.67 8.97 13.04 4.21
E.2 Correlation between prior and PPL
@@
Appendix F Details on experiments
F.1 Scores for each benchmark
@@
ARC easy BIG-bench wikidata TriviaQA COPA OBQA PIQA HellaSwag LAMBADA Winograd Winogrande
@@
BIG-bench algorithms BIG-bench dyck lan- guages BIG-bench elementary math QA BIG-bench operators GSM8K SVAMP LSAT-LR LSAT-RC SAT-English CoQA
@@
F.2 Justification of experiment scale
F.2.1 Regarding training token duration
To ensure experimental rigor despite our resource constraints, we followed the experimental setup of a representative paper that we conceptually inherit Ankner et al. (2024). According to this study, the performance gap between baselines is established in the early phases of training. Ankner et al. (2024) reports downstream task performance up to 25 or 50B token duration, and it shows that the performance gaps between baselines observed at 5B token duration remain largely consistent throughout the last stage (Figure 1 of Ankner et al. (2024)). Our work leverages this observation, and this also aligns with our experiment in Figure 7.
@@ F.2.1 Regarding training token duration
Particularly in Ankner et al. (2024), all models are trained with 2× data repetition, and the 5B-token duration in Figure 1 of Ankner et al. (2024) corresponds to 2.5B tokens of data repeated twice. Our experimental setup also uses 3B tokens repeated twice (total 6B), which makes it directly comparable to the 5B-token point in Figure 1 of Ankner et al. (2024).
These trends are consistently observed in other works. li2024datacomp also demonstrates that the ranking of model performance is largely determined early in training and remains consistent. In addition, Figure 1 of diao2025climb clearly shows that the initial performance gaps are established in the early phases.
F.2.2 Regarding the model size
Many representative works on data filtering employ relatively small models around 1.5B parameters, assuming that such a scale is sufficient for reliable evaluation. For example, Ankner et al. (2024) uses 1B and 2B models, while Diao et al. (2025) employs 350M and 1B models. In addition, recent representative works in this topic (Li et al., 2024b; Penedo et al., 2024; Longpre et al., 2024; Xie et al., 2023b) all conduct experiments with models smaller than 2B. This contrasts with studies focusing on model architectures (e.g., LLaMA, DeepSeek) that compete in complex reasoning performance, which typically require pretraining of large-scale models (e.g., 7B and 70B). A detailed rationale of why small model settings can still maintain experimental rigor in data filtering, is provided below.
F.3 Indirect evaluation on the complex reasoning task
Certain abilities, like complex reasoning skills, may exhibit emergence in the later stages of training or in larger models. This may be why the reviewers expect experiments at a larger scale. We will provide an explanation and an indirect evaluation regarding this.
F.3.1 Justification
Our method is designed to filter noisy text data, which consists of broken sentence structures. If we can assume noise as "data that contains no interpretable information", then learning such data does not contribute to the emergence of complex reasoning. This is because the emergence of complex intelligence arises from the integration of information within the model (Cleveland, 1982), but noise contains no information to contribute to this process. See noise cases filtered by prior-filter in Appendix F; they will remain useless no matter how much the model learns it or how large model learns it.
If we can assume that learning more information (i.e., from corpora with less noise) leads to better complex reasoning skills, the amount of learned information can be evaluated at a smaller scale and with more straightforward benchmarks, which are sufficiently covered in our experiments.
F.3.2 Indirect evaluation
Recently, some research views that reasoning skill is learned directly from the dataset as a part of pattern memorization, rather than emerging as a mere intellectual invention (Wu et al., 2024; Mirzadeh et al., 2024). From this view, if the training dataset contains more data that is similar to the targeted complex task benchmark, the model’s performance on that target benchmark would increase accordingly. This is because the model can utilize a larger number of direct templates to solve the task, in a manner of pattern memorization.
From this concept, we mixed the targeted benchmark dataset (e.g., MMLU-pro) into the Dolma corpus and measured how much of this benchmark data is maintained through each filtering method. This simulates a scenario where the training dataset contains data similar to the target benchmark. Therefore, if more benchmark data remain after filtering, we can assume there are higher likelihood that such data (i.e., similar to the targeted benchmark) is preserved in a larger-scale setting, leading to better benchmark performance. Although we can not directly evaluate MMLU-pro performance with a small LM (usually under 7B), this may offer an indirect way to estimate the potential impact of filters on complex task-solving capability.
Table 12: The remaining portion of the MMLU-Pro dataset for each filtering method and threshold.
Threshold (<) ±0.2 ±0.25 ±0.3 ±0.4 ±0.45
PPL-based 0.222 0.292 0.436 0.740 0.987
Prior-based 0.510 0.629 0.743 0.931 0.990
Random 0.400 0.500 0.600 0.800 0.900
As a result, we find that the prior-based filter achieves significantly better recognition performance on MMLU-Pro data compared to PPL and random selection. Based on this observation, it is difficult to argue that a model trained with PPL-based filter would achieve a better score than a prior-based filter on MMLU-Pro.
We analyze that this might be due to our method’s better preservation of “rare but valuable data,” as discussed in Appendix B.1. MMLU-Pro is known to contain a lot of logical statements, which likely constitute only a small portion of web corpora (e.g., Dolma). As prior-filter is more sensitive to the rare group (§D), it might have a higher chance of capturing such data.
Appendix G Details on Benchmarks
@@
Jha et al. (2023) also use the MosaicML evaluation gauntlet to perform evaluations in their work. As such, with explicit permission from the authors, we reproduce their text describing the tasks and task categories in the evaluation gauntlet. The following is from Section D of their paper:
@@
  • ARC easy: 2,376 easy four-choice multiple choice science questions drawn from grade 3-9 science exams. (Clark et al., 2018)
  • BIG-bench wikidata: 20,321 questions regarding factual information pulled from Wikipedia. (Srivastava et al., 2023)
  • TriviaQA: 3,000 question-answering dataset; clipped all answers to be at most 10 tokens long to improve speed. (Joshi et al., 2017)
@@
  • COPA: 100 cause/effect multiple choice questions. (Roemmele et al., 2011)
  • OBQA (OpenBook QA): 500 four-choice multiple choice questions that rely on basic physical and scientific intuition about common objects and entities. (Mihaylov et al., 2018)
  • PIQA: 1,838 commonsense physical intuition 2-choice multiple choice questions. (Bisk et al., 2020)
@@
  • HellaSwag: 10,042 multiple choice scenarios in which the model is prompted with a scenario and choose the most likely conclusion to the scenario from four possible options. (Zellers et al., 2019)
  • LAMBADA: 6,153 passages take from books - we use the formatting adopted by OpenAI’s version. (Paperno et al., 2016)
  • Winograd Schema Challenge: 273 scenarios in which the model must use semantics to correctly resolve the anaphora in a sentence. (Levesque et al., 2012)
  • Winogrande: 1,267 scenarios in which two possible beginnings of a sentence are presented along with a single ending. (Sakaguchi et al., 2021)
@@
  • BIG-bench algorithms: 1,320 multiple choice questions. (Srivastava et al., 2023)
  • BIG-bench dyck languages: 1000 complete-the-sequence questions. (Srivastava et al., 2023)
  • BIG-bench elementary math QA: 38,160 four-choice multiple choice arithmetic word problems. (Srivastava et al., 2023)
  • BIG-bench operators: 210 questions involving mathematical operators. (Srivastava et al., 2023)
  • GSM8K: 1,319 short, free-response grade school-level arithmetic word problems with simple numerical solutions. (Cobbe et al., 2021)
  • SVAMP: 300 short, free-response grade school-level arithmetic word problems with simple numerical solutions. (Patel et al., 2021)
@@
  • LSAT-LR: 510 passage-based four choice multiple choice questions. (Zhong et al., 2023)
  • LSAT-RC: 268 passage-based four choice multiple choice questions. (Zhong et al., 2023)
  • SAT-English: 206 passage-based four choice multiple choice questions. (Zhong et al., 2023)
  • CoQA: 7,983 passage-based short free response questions. (Reddy et al., 2019)
Appendix H Usage of AI assistants
@@
Appendix I Reproducibility statement
Within the abstract, we included a URL to an anonymous GitHub repository where our code is made available. This repository not only contains the full implementation but also offers detailed guidelines for installation, along with step-by-step instructions on how to perform both training and testing.
Appendix J Filtered cases
@@ Appendix I Reproducibility statement
J.1 Prior stds($\sigma_{\mathtt{d}}\$) based
@@
J.2 Prior mean($\mu_{\mathtt{d}}\$)-based